OpenStudy (anonymous):

The sum of two integers is 15. Eight times the smaller integer is one less than three times the lager integer. Find the integers. This problem should not use system of equations. I see how to set the problem up using system of equations. How do you set the problem up using one equation similar to a consecutive equation?

jhonyy9 (jhonyy9):

try writing what you know x +y = 15 8x -1=3y so hope i have understood it right your exercise

OpenStudy (loser66):

@jhonyy9 he said that not use system of equation

OpenStudy (anonymous):

Yes, the problem should not use system of equation but one equation similar to a consecutive

OpenStudy (anonymous):

Yes, I was looking at it like this 8x + 3x -1 = 15 8x = small' 3x-1 = large but it does not come out right

OpenStudy (anonymous):

8x +8x-1=15 is not correct The integers should be 4 and 11

OpenStudy (anonymous):

If you set this problem up using system of equation, you get the integer 4 and 11 and that is the answer in the book but they do not set the problem up using system of equation. They also don't show how the problem was setup.

OpenStudy (anonymous):

They are right. http://www.wolframalpha.com/input/?i=x+%2B+y+%3D+15+and+8x+-+3y+%3D+-1

OpenStudy (anonymous):

The link is the problem set up as a system of equations.

OpenStudy (loser66):

yea, we have to solve by system of equations.

OpenStudy (anonymous):

No, there is a way to set it up not using systems of equations. The chapter the problem is in does not even touch on system of equation and they point to a consecutive example but it is not like this. The example is not. This makes perfect sense to me a system of equation but how to set it up with out using system of equation. That is what is confusing me. I just don't see it.

OpenStudy (anonymous):

Their example they point to is: One number is four more than another number. The sum of the two numbers is sixty six. Find the two numbers I get this one. n + (n+4) = 66 Makes perfect sense when reading it but the other one is just not clicking.

OpenStudy (loser66):

@ParthKohli

Parth (parthkohli):

oh... yeah

OpenStudy (anonymous):

Yes, it should not be a system of equations.

Parth (parthkohli):

See, if the first integer is \(x\) then the larger one is represented by \(\dfrac{8x + 1}{3}\)

OpenStudy (anonymous):

Yes he got the smaller one and I was trying to see how he come to that conclusion

Parth (parthkohli):

Huh?\[8x = 3y - 1 \Rightarrow \dfrac{8x + 1}{3} = y\]

OpenStudy (loser66):

oh, ya,

OpenStudy (loser66):

thanks ParthKohli

OpenStudy (anonymous):

This is not like a consecutive problem, which the book points out. I see an undstand what Parth did.

Parth (parthkohli):

That's it. So\[x + \dfrac{8x + 1}{3}\]

Parth (parthkohli):

\[= 15\]

OpenStudy (loser66):

@MDoodler he used consecutive number to construct 8x+1/3

OpenStudy (loser66):

However, it's kind of break down the system of equation. hihihi...

OpenStudy (anonymous):

so he 8x = 3y - 1 , which I see that in the word problem and then move -1 to the 8x side and divided by 3 to get y alone. I see this but why the books through the whole chapter 2 not give a single problem like this. Just makes me think they are doing something else. Anyways thanks for the help. I might post later to see what others thoughts are on the subject.

Parth (parthkohli):

Exactly, a cheated a little...